### v1

@@ VII-D Fine-tuned LLM Prediction
Fine-tuning pre-trained LLM models presents a significant advancement for time-series prediction, offering a powerful alternative to traditional prediction approaches [190, 191]. General-domain LLM models, initially pre-trained on extensive linguistic data, can be fine-tuned to capture the unique temporal patterns inherent in time-series data. This process equips LLMs with the ability to effectively prediction in domains where data scarcity or specificity presents challenges to conventional deep learning models. In the pursuit of efficiency and practicality, most recent works have shifted towards parameter-efficient fine-tuning methods like Low-Rank Adaptation (LORA) [192] and Layer Normalization Tuning (LNT) [193]. In particular, LORA adapts pre-trained models to new tasks by modifying the weight matrices of the model’s layers. Given a weight matrix $W\in\mathbb{R}^{d\times m}$ in a pre-trained model, LORA fine-tuning introduces two low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times m}$, where $r$ is the rank and $r\ll\min(d,m)$. The weight matrix $W$ is updated as:
@@
The matrices $A$ and $B$ are the parameters learned during fine-tuning while the original weights $W$ are kept frozen. This results in a model that is fine-tuned for the task at hand with only a small increase in the number of parameters. Many works of fine-tuning LLMs proposed applying the technique to the query (Q) and value (V) matrices in attention layers, showing notable results without extending it to all parameters within the attention or feed-forward layers. In the context of time-series prediction, however, this selective fine-tuning may require adjustment. As shown in Fig. 19, LLM4TS [190] applies LORA fine-tuning to the query (Q) and key (K), achieving state-of-the-art performance. Using LORA allows for retaining the general capabilities of the LLM while imbuing it with domain-specific knowledge, ensuring that the time-series prediction model is both specialized and robust.
@@
On the other hand, LNT offers a focused approach to adapt pre-existing parameters in transformer blocks to specific tasks. LNT specifically targets the affine transformation parameters within the layer normalization components of a transformer model. These parameters, such as scale and shift, originally set to ensure standardized input distribution across network layers, become trainable to allow the model to retain its learned representations while fine-tuning the time-series prediction. As shown in Fig. 19, LLM4TS [190] employs both LNT and LORA fine-tuning for the query and key. A similar strategy can be found in [191], which freezes all attention and feed-forward layers, and only fully fine-tunes the embedding layers and applies the LNT. Incorporating LNT in the fine-tuning process, in the context of adapting pre-trained LLM models for time-series prediction, provides a mechanism for the model to adjust its internal normalization to better fit the dynamics and scale of the time-series data.
These parameter-efficient fine-tuning methods, such as LoRA and LNT, are crucial for the practical deployment of LLM models in telecom. Lin et al. claims that applying LoRA to GPT-3 can reduce the number of trainable parameters from 175.2 billion to 37.7 million[16], and combining LoRA with federated split learning can significantly reduce computing and communication latency at the mobile edge.
@@ VII-E Multi-modal LLM for Telecom Prediction
Pre-training foundation models Leverages pre-trained models on diverse time series datasets to capture general temporal patterns. It means training LLM models from scratch specifically for prediction purposes. Requires a large corpus of time-series data for initial pre-training, but collecting these datasets may be difficult in telecom; fine-tuning may be needed for specific telecom tasks. The model has zero-shot prediction capabilities, and quickly adapts to new tasks with minimal fine-tuning; captures a wide range of temporal dynamics. A prediction foundation model for telecom can handle various short-term or long-term prediction tasks, such as traffic load prediction, CSI prediction, user number estimation, and so on.
Frozen pre-trained LLM for prediction Using pre-trained LLM models without fine-tuning their parameters, including prompting-based and preprocessing-based methods. Tokenization and embedding of time series data; For prompt-based methods, the prompt format must be carefully designed; No fine-tuning for LLM models is required. Low computational cost and design complexity; Leveraging the generalization capabilities of a pre-trained LLM directly. This technique is particularly useful for short-term prediction, such as short-term traffic load and network performance prediction. The low computational cost can also adapt to network edge, even mobile applications.
Fine-tuned LLM prediction Adapts a pre-trained LLM to telecom-specific prediction tasks through fine-tuning techniques such as LoRA and LNT. It requires time series data for fine-tuning; may require parameter-efficient fine-tuning like LoRA and LNT to improve the efficiency. Fine-tuning can incorporate telecom domain knowledge into LLM models, improving the accuracy and specificity of telecom tasks. Fine-tuning LLMs can better adapt to specific tasks in telecom, e.g., collecting specific datasets to fine-tune an LLM for user localization. It is more flexible than pre-trained models from scratch, and more reliable than pure prompting-based methods.
Multi-modality prediction Using LLM models to jointly consider multi-modal environment information, e.g., tabular data, text, and image, aiming to provide more accurate prediction results. It requires multi-modal input and proper prompt to predict desired output; LLM models can be specifically pre-trained/fine-tuned to further improve the performance and generalization capabilities. Multi-modal can take advantage of inputs with various modalities. With these comprehensive inputs, LLMs can better predict the network dynamics than existing methods. Sensing is an important part of 6G networks, and multi-modal sensing can provide more comprehensive input for LLM models, producing more accurate prediction results by jointly considering various inputs, e.g., more accurate traffic load prediction and beam steering.
@@ VIII-A1 Telecom-domain LLM training
Previous sections have shown the importance of building telecom-specific LLMs, e.g., telecom-domain question answering[100], telecom troubleshooting[23], and standard specification classification[13]. Despite the great potential, training an LLM model specifically for telecom presents unique challenges due to the complex nature of large-scale communication networks. In the following, we will analyze this challenge in detail.
@@
Sufficient domain-specific datasets are prerequisites for training a telecom LLM. Unlike general-domain LLM models, which can leverage large-scale text corpora, obtaining a sizable dataset exclusively focused on communication networks can be challenging. Existing studies usually focus on one specific task and then build the corresponding dataset, e.g., the trouble report dataset [23], 3GPP specification dataset [13], and telecom question answering dataset [100]. However, these datasets are usually small-scale and task-specific, and a comprehensive large-scale dataset should include network-related documents, standard specifications, protocols, textbooks, research papers, and other relevant sources. Maatouk et al. started the exploration in [110] by building a dataset with 10000 telecom-related questions and answers, including around 25000 pages and 6 million words. More efforts are needed to provide more comprehensive and diverse datasets for telecom LLM training.
@@
Meanwhile, it is worth noting that telecom networks involve a large number of various concepts such as network protocols, routing algorithms, network topologies, network security, etc. Therefore, teaching an LLM to comprehend and reason about these complex concepts requires a robust training strategy. An effective approach is to pre-train the LLM on a large-scale general language corpus and then fine-tune it on specific communication network datasets, e.g., datasets for the BS services, historical datasets for prediction, or datasets for edge computing-related tasks. In addition, balancing model size and performance is crucial. LLM models trained on large-scale datasets tend to be computationally expensive and memory-intensive. Appropriate model size can reduce the burden on energy and computation resources during pre-training and fine-tuning phases. In addition, balancing model size and performance is crucial to ensure practical usability, especially considering scenarios with limited computational capacity such as vehicles and mobile phones. Therefore, techniques like model compression, knowledge distillation, or utilizing specialized hardware accelerators can be explored to reduce the model’s size and enhance its efficiency without compromising its understanding of communication networks.
In summary, training an LLM specifically for telecom is of great importance. However, multiple challenges must be addressed, including domain-specific training data, complex networking concepts, and balancing model size and performance. Overcoming these challenges will enable the development of a robust LLM capable of effectively applying networking domain knowledge to various telecom tasks.
@@ VIII-A2 Practical LLM deployment in telecom
To leverage the benefits of LLM techniques, the models should be properly deployed in telecom networks. Specifically, LLM models can be deployed at different levels, including central cloud, network edge, or user devices. We have introduced the features of each approach in Section III-F. However, the related studies are still in very early stages, and the proposed schemes mainly focus on system-level design and definitions. The following will discuss the key challenges and difficulties for practical LLM deployment in telecom networks.
Firstly, many real-world wireless applications have stringent requirements for service delay, e.g., autonomous driving and robotic control. With such time constraints, using a central cloud-based LLM to process these latency-critical tasks can be inappropriate, since the task uploading and solution downloading may increase the service delay. Additionally, if the task involves image and video processing, the uploading and downloading process will significantly increase the latency, especially considering the limited backhaul capacity. For instance, the image classification tasks introduced in Section V-D require rapid responses for signal blockage prediction and autonomous driving[25] [146], and processing these requirements on cloud can be impractical due to high service latency. In addition, the LLM inference time will also contribute to system latency, ranging from 0.58 to 90 seconds. Therefore, the service time should be very carefully evaluated before using a LLM for latency-critical applications.
Network edge provides an efficient approach for computational task processing, and edge intelligence has become an appealing direction to deploy ML algorithms in telecom networks, especially for distributed algorithms such as federated learning. However, network edge servers have limited computational or storage capacity, and LLMs are usually computationally intensive with large model sizes, which may prevent edge-LLM deployment. By contrast, hybrid deployment approaches, combining central cloud, edge, and user device deployments, can provide a balance between scalability, low latency, and privacy. For example, a hierarchical architecture can be used, where various LLMs are deployed at the central cloud for resource-intensive tasks, at the network edge for latency-sensitive tasks, and on user devices for personalized or offline tasks. However, coordinating these different deployment levels can be challenging and requires efficient protocols and algorithms. Xu et al. proposed a hybrid deployment approach in [17], but this is still an open challenge for the communication society. Deploying LLM models at different levels, including the central cloud, network edge, or user devices, offers unique opportunities and challenges in telecom applications. Each deployment methods has its considerations regarding latency and computational resources. Hybrid deployment can be an ideal solution, leveraging the advantages of multiple deployment levels, but the coordination may lead to potential issues.
@@ VIII-A3 Prompt engineering for telecom applications
Prompt engineering is a crucial aspect of utilizing LLM techniques effectively, as it plays a significant role in guiding the model’s behaviour and generating desired outputs. However, designing prompts for telecom applications presents unique challenges due to the domain complexity. This subsection will discuss the challenges of prompt engineering for telecom networks.
Telecom networks encompass a wide range of concepts, protocols, and technologies, making it challenging to distill the necessary information into a concise prompt. The diverse nature of the domain requires a deep understanding of networking principles and the ability to capture specific nuances related to network architectures, protocols, performance optimization, and security. To design effective prompts, researchers must identify the most relevant components and provide concise yet comprehensive instructions to LLMs.
Meanwhile, prompt designs should strike a balance between being specific enough to guide the LLM in generating accurate and contextually appropriate responses, while also remaining general enough to handle a wide range of network-related queries or tasks. Achieving this balance is crucial as overly specific prompts may limit the model’s ability to generalize, while a general prompt may lead to irrelevant responses. Fine-tuning the prompt language and incorporating domain-specific keywords or phrases can help achieve the desired balance for telecom applications.
@@
To summarize, prompt design of LLMs for telecom applications poses a significant challenge due to the intricate and constantly evolving nature of the domain. Crafting effective prompts necessitates a profound comprehension of networking principles, the capacity to strike a balance between specificity and generality, and an awareness of contextual factors. Collaboration and input from researchers and industry experts play a crucial role in developing robust prompt engineering strategies that empower LLM agents to deliver precise, pertinent, and impartial responses in telecom applications.
@@ VIII-B Future Directions
This subsection will identify possible future directions for LLM-empowered telecom.
VIII-B1 LLM-enabled planning in telecom

### v2

@@ VII-D Fine-tuned LLM Prediction
Fine-tuning pre-trained LLMs presents a significant advancement for time-series prediction, offering a powerful alternative to traditional prediction approaches [205, 206]. General-domain LLMs, initially pre-trained on extensive linguistic data, can be fine-tuned to capture the unique temporal patterns inherent in time-series data. This process equips LLMs with the ability to effectively prediction in domains where data scarcity or specificity presents challenges to conventional deep learning models. In the pursuit of efficiency and practicality, most recent works have shifted towards parameter-efficient fine-tuning methods like Low-Rank Adaptation (LORA) [207] and Layer Normalization Tuning (LNT) [208]. In particular, LORA adapts pre-trained models to new tasks by modifying the weight matrices of the model’s layers. Given a weight matrix $W\in\mathbb{R}^{d\times m}$ in a pre-trained model, LORA fine-tuning introduces two low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times m}$, where $r$ is the rank and $r\ll\min(d,m)$. The weight matrix $W$ is updated as:
@@
The matrices $A$ and $B$ are the parameters learned during fine-tuning while the original weights $W$ are kept frozen. This results in a model that is fine-tuned for the task at hand with only a small increase in the number of parameters. Many works of fine-tuning LLMs proposed applying the technique to the query (Q) and value (V) matrices in attention layers, showing notable results without extending it to all parameters within the attention or feed-forward layers. In the context of time-series prediction, however, this selective fine-tuning may require adjustment. As shown in Fig. 20, LLM4TS [205] applies LORA fine-tuning to the query (Q) and key (K), achieving state-of-the-art performance. It augments the pre-trained model with additional trainable components and thus incorporates changes to the model’s weights, unlike the soft prompting to modify the inputs. Using LORA allows for retaining the general capabilities of the LLM while imbuing it with domain-specific knowledge, ensuring that the time-series prediction model is both specialized and robust.
@@
On the other hand, LNT offers a focused approach to adapt pre-existing parameters in transformer blocks to specific tasks. LNT specifically targets the affine transformation parameters within the layer normalization components of a transformer model. These parameters, such as scale and shift, originally set to ensure standardized input distribution across network layers, become trainable to allow the model to retain its learned representations while fine-tuning the time-series prediction. As shown in Fig. 20, LLM4TS [205] employs both LNT and LORA fine-tuning for the query and key. A similar strategy can be found in [206], which freezes all attention and feed-forward layers, and only fully fine-tunes the embedding layers and applies the LNT. Incorporating LNT in the fine-tuning process, in the context of adapting pre-trained LLMs for time-series prediction, provides a mechanism for the model to adjust its internal normalization to better fit the dynamics and scale of the time-series data.
These parameter-efficient fine-tuning methods, such as LoRA and LNT, are crucial for the practical deployment of LLMs in telecom. Lin et al. claims that applying LoRA to GPT-3 can reduce the number of trainable parameters from 175.2 billion to 37.7 million[16], and combining LoRA with federated split learning can significantly reduce computing and communication latency at the mobile edge.
@@ VII-E Multi-modal LLM for Telecom Prediction
Pre-training foundation models Leverages pre-trained models on diverse time series datasets to capture general temporal patterns. It means training LLMs from scratch specifically for prediction purposes. Requires a large corpus of time-series data for initial pre-training, but collecting these datasets may be difficult in telecom; fine-tuning may be needed for specific telecom tasks. The model has zero-shot prediction capabilities, and quickly adapts to new tasks with minimal fine-tuning; captures a wide range of temporal dynamics. A prediction foundation model for telecom can handle various short-term or long-term prediction tasks, such as traffic load prediction, CSI prediction, user number estimation, and so on.
Frozen pre-trained LLM for prediction Using pre-trained LLMs without fine-tuning their parameters, including prompting-based and preprocessing-based methods. Tokenization and embedding of time series data; For prompt-based methods, the prompt format must be carefully designed; No fine-tuning for LLMs is required. Low computational cost and design complexity; Leveraging the generalization capabilities of a pre-trained LLM directly. This technique is particularly useful for short-term prediction, such as short-term traffic load and network performance prediction. The low computational cost can also adapt to network edge, even mobile applications.
Fine-tuned LLM prediction Adapts a pre-trained LLM to telecom-specific prediction tasks through fine-tuning techniques such as LoRA and LNT. It requires time series data for fine-tuning; may require parameter-efficient fine-tuning like LoRA and LNT to improve the efficiency. Fine-tuning can incorporate telecom domain knowledge into LLMs, improving the accuracy and specificity of telecom tasks. Fine-tuning LLMs can better adapt to specific tasks in telecom, e.g., collecting specific datasets to fine-tune an LLM for user localization. It is more flexible than pre-trained models from scratch, and more reliable than pure prompting-based methods.
Multi-modality prediction Using LLMs to jointly consider multi-modal environment information, e.g., tabular data, text, and image, aiming to provide more accurate prediction results. It requires multi-modal input and proper prompt to predict desired output; LLMs can be specifically pre-trained/fine-tuned to further improve the performance and generalization capabilities. Multi-modal can take advantage of inputs with various modalities. With these comprehensive inputs, LLMs can better predict the network dynamics than existing methods. Sensing is an important part of 6G networks, and multi-modal sensing can provide more comprehensive input for LLMs, producing more accurate prediction results by jointly considering various inputs, e.g., more accurate traffic load prediction and beam steering.
@@ VIII Challenges and Future Directions of LLM-empowered telecom
TABLE XII: Summary of Telecom datasets for LLM
Dataset Task Document size Question size Open-source
5GSum[219] Summarization 713 articles N/A Yes
Tspec-LLM[220] Summarization 30,137 documents 100 questions Yes
TeleQnA[125] Question answering N/A 10,000 questions Yes
NetEval[221] Question answering N/A 5,732 questions Yes
TeleQuAD[115] Question answering N/A 2,021 questions No
StandardsQA[222] Question answering N/A 2,400 questions No
ORAN-Bench-13K[223] Question answering 116 documents 13,952 questions Yes
5GSC[219] Sentence classification 2,401 sentences N/A Yes
@@ VIII-A1 Telecom-domain LLM training
Previous sections have shown the importance of building telecom-specific LLMs, e.g., telecom-domain question answering[115], telecom troubleshooting[23], and standard specification classification[13]. Despite the great potential, training an LLM specifically for telecom presents unique challenges due to the complex nature of telecom networks. In the following, we will analyze this challenge in detail.
@@
Sufficient telecom-related datasets are prerequisites for training a telecom LLM. Unlike general-domain LLMs, which can leverage large-scale text corpora on the internet, obtaining a sizable dataset exclusively focused on communication networks can be challenging. Existing studies usually focus on one specific task and then build the corresponding dataset, e.g., the trouble report dataset [23], 3GPP specification dataset [13], and telecom question answering dataset [115]. However, these datasets are usually small-scale and task-specific, and a comprehensive large-scale dataset should include network-related documents, standard specifications, protocols, textbooks, research papers, and other relevant sources. Maatouk et al. started the exploration in [125] by building a dataset with 10000 telecom-related questions and answers, including around 25000 pages and 6 million words. More efforts are needed to provide more comprehensive and diverse datasets for telecom LLM training.
@@
Meanwhile, it is worth noting that telecom networks involve a large number of various concepts such as network protocols, routing algorithms, network topologies, network security, etc. Therefore, teaching an LLM to comprehend and reason about these complex concepts requires a robust training strategy. An effective approach is to pre-train the LLM on a large-scale general language corpus and then fine-tune it on specific communication network datasets, e.g., datasets for the BS services, historical datasets for prediction, or datasets for edge computing-related tasks. In addition, balancing model size and performance is crucial. LLMs trained on large-scale datasets tend to be computationally expensive and memory-intensive. Appropriate model size can reduce the burden on energy and computation resources during pre-training and fine-tuning phases. In addition, balancing model size and performance is crucial to ensure practical usability, especially considering scenarios with limited computational capacity such as vehicles and mobile phones. Therefore, techniques like model compression, knowledge distillation, or utilizing specialized hardware accelerators can be explored to reduce the model’s size and enhance its efficiency without compromising its understanding of communication networks.
The above analyses show that obtaining domain-specific datasets is one of the main bottlenecks of training telecom LLMs. Several datasets have been released recently to address this challenge. As shown in Table XII, there exist different types of datasets, focusing on tasks including text summarization, question answering, and sentence classification. Those datasets are mainly extracted from telecom-related documents. By integrating these datasets, LLMs can be trained to understand and respond to customer queries more effectively, predict and mitigate network issues, and identify fraudulent activities with higher accuracy, ultimately leading to improved operational efficiency and customer satisfaction in the telecom sector.
@@ VIII-A2 Practical LLM deployment in telecom
To leverage the benefits of LLM techniques, the models should be properly deployed in telecom networks. Specifically, LLMs can be deployed at different levels, including central cloud, network edge, or user devices. We have introduced the features of each approach in Section III-F. However, the related studies are still in very early stages, and the proposed schemes mainly focus on system-level design and definitions. The following will discuss the key challenges and difficulties for practical LLM deployment in telecom networks.
Firstly, many real-world wireless applications have stringent requirements for service delay, e.g., autonomous driving and robotic control. With such time constraints, using a central cloud-based LLM to process these latency-critical tasks can be inappropriate, since the task uploading and solution downloading may increase the service delay. Additionally, if the task involves image and video processing, the uploading and downloading process will significantly increase the latency, especially considering the limited backhaul capacity. For instance, the image classification tasks introduced in Section V-D require rapid responses for signal blockage prediction and autonomous driving[25] [161], and processing these requirements on cloud can be impractical due to high service latency. In addition, the LLM inference time will also contribute to system latency, ranging from 0.58 to 90 seconds. Therefore, the service time should be very carefully evaluated before using a LLM for latency-critical applications. Network edge provides an efficient approach for computational task processing, and edge intelligence has become an appealing direction to deploy ML algorithms in telecom networks. However, network edge servers have limited computational or storage capacity, and LLMs are usually computationally intensive with large model sizes, which may prevent edge-LLM deployment.
To this end, hybrid deployment can be an ideal solution by combining central cloud, edge, and user device deployments, providing a balance between scalability, low latency, and privacy. Deploying LLMs at different levels, including the central cloud, network edge, or user devices, offers unique opportunities and challenges in telecom applications. For example, large-scale LLMs such as GPT-4 and LLama3-70b are deployed at the central cloud to handle tasks with high-quality requirements on the generated content. Meanwhile, small-scale LLMs are implemented at the network edge or even on user devices for latency-sensitive tasks. However, coordinating LLMs at different levels can be challenging and requires dedicated designs. For example, how to select appropriate LLMs for diverse user tasks such as lower latency and price, higher generation quality, or multi-modal tasks.
@@ VIII-A3 Prompt engineering for telecom applications
Prompt engineering is a crucial aspect of utilizing LLM techniques effectively, as it plays a significant role in guiding the model’s behaviour and generating desired outputs. Specifically, prompt engineering refers to the process of structuring an instruction that can be interpreted and understood by generative AI models, and then produce the desired output. For instance, few-shot learning can be considered one of the prompt engineering approaches, in which LLMs can learn from examples and demonstrations to improve their performance on target tasks. Compared with pre-training or fine-tuning an LLM, prompting has a much lower cost on computational resources. In particular, it relies on LLM’s inference capabilities, indicating the most straightforward and efficient approach to use LLMs. The high efficiency of prompting techniques aligns well with many telecom applications, which usually require fast responses to network dynamics, e.g., changing channel conditions, user numbers, network traffic level, etc. However, designing prompts for telecom applications presents unique challenges due to the domain complexity.
Firstly, telecom networks encompass a wide range of concepts, protocols, and technologies, making it challenging to distill the necessary information into a concise prompt. The diverse nature of the domain requires a deep understanding of networking principles and the ability to capture specific nuances related to network architectures, protocols, performance optimization, and security. To design effective prompts, researchers must identify the most relevant components and provide concise yet comprehensive instructions to LLMs. Meanwhile, prompt designs should strike a balance between being specific enough to guide the LLM in generating accurate and contextually appropriate responses, while also remaining general enough to handle a wide range of network-related queries or tasks. Achieving this balance is crucial as overly specific prompts may limit the model’s ability to generalize, while a general prompt may lead to irrelevant responses.
@@
To summarize, prompt design of LLMs for telecom applications poses a significant challenge due to the intricate and constantly evolving nature of the domain. Crafting effective prompts necessitates a profound comprehension of networking principles, the capacity to strike a balance between specificity and generality, and an awareness of contextual factors. To this end, a practical solution is to publicize some standard prompting templates. They will provide fundamental suggestions for prompt designs of each kind of task. For instance, network optimization-related tasks should specify the optimization objective and control variables, providing feedback or examples for previous selections. In addition, researchers and industry experts play a crucial role in developing these standards, since the template design requires professional knowledge and understanding of telecom networks. Then these templates will be able to deliver precise, pertinent, and impartial responses in telecom applications.
@@ VIII-B Future Directions
VIII-B1 Multi-modal LLMs for telecom
Multi-modality is a crucial direction for LLM development, indicating seamless integration of information with various modalities such as text, image, audio, video, etc. Such a capability may serve many applications in future telecom networks. For instance, sensing has become a critical pillar for envisioned 6G networks, and multi-modal LLMs can utilize 3D multi-modal data, e.g., text, satellite or street camera images, 3D LiDAR maps and videos, to provide a holographic understanding for wireless signal transmission environment [30]. A specific example of mmWave/THz beamforming has been discussed in Section VII-E, and it aims to predict signal transmission blockage by using multi-modal input such as image and video. Meanwhile, with multi-modal information of the 3D environment, we can also have better CSI estimation results, which is critical for signal transmission and network management [30]. In addition, Xu et. al also introduced an example in [17], which utilizes LLMs to generate a traffic accident report by using videos collected by vehicles. This video-to-text generation can also be used to analyze the videos collected by UAVs to describe the wireless signal transmission environments.
VIII-B2 LLM-enabled planning in telecom